At a circus, 135 clowns are getting into cars. Blue cars hold 20 clowns and red cars only hold 15. How many of each are there?
Answer by stanbon(75887)

At a circus, 135 clowns are getting into cars. Blue cars hold 20 clowns and red cars only hold 15. How many of each are there?
---
20b + 15r = 135
20b = 135-15r
4b = 27 - 3r
b = (27-3r)/4
---
b = (3/4)(9-r)
-----
Graph and look for whole number solutions.
-----------------------------------------------
b = 6 when r = 1
---
Answers:
6 blue cars and 1 red car
OR
3 blue cars and 5 red cars
Or
0 blue cars and 9 red cars
===================================
Cheers,
Stan H.